A swimming pool is in the shape of a rectangle. The length of the pool is 15 feet more than the width. A concrete deck 6 feet wide is added around the pool, and the perimeter of the deck is 178 feet. Find the dimension of the swimming pool.
Answer by ankor@dixie-net.com(22740)

A swimming pool is in the shape of a rectangle.
The length of the pool is 15 feet more than the width.
A concrete deck 6 feet wide is added around the pool, and the perimeter of the deck is 178 feet.
Find the dimension of the swimming pool.
:
Let x = the width of the pool
"The length of the pool is 15 feet more than the width. "; therefore
(x+15) = the length
:
The 6' deck goes all the way around the pool therefore we add 12' to each dimension of the pool
x + 12 = the overall width
x + 15 + 12 = x+27 is overall length
:
The perimeter
2(x+27) + 2(x+12) = 178
divide eq by 2
x + 27 + x + 12 = 178/2
2x + 39 = 89
2x = 89 - 39
2x = 50
x = 25' is the width of the pool
and
25 + 15 + 40' is length
;
:
Check this using the perimeter
2(52) + 2(37) =
104 + 74 = 178